30x² - 20x
If we factored it incorrectly, say, we only took out a 10.
10(3x² - 2x) This is not factored correctly but if we check it, it will go back to the original problem.
30x² - 20x.
How could you check to make sure you took out the GCF? Is there a way to find the GCF between 30 and 20? Then you could do the variables separate.

Answer by mananth(16949)

30x² - 20x
the factors for
30 = 2*3*5
20= 2*2*5
..
2& 5 are common in both
2*5 =10 is the GCF
..
10x(3x-2)